the future value of $7500 in 3 years with simole interest rate of 5% is?
Answer #1

The total interest earned in 3 years = Present value * rate of interest* number of years = $7,500 * 5%* 3 = $7,500 * 5/100 * 3 = $75*15= $1,125.

So, the future value = principal + interest earned = $7,500 + $1,125 = $8,625.